OpenStudy (anonymous):

how do you find the slope and y intercept of x-y=-2

OpenStudy (abbles):

First you need to put it into y-intercept form (y = mx + b) by isolating the y and the x is the slope and the b is the y-intercept.

OpenStudy (abbles):

To isolate the y, first subtract x from both sides. x - y = -2 -y = -x - 2 Now you can multiply the whole equation by -1 to turn everything positive. y = x - 2 Now you can tell that the slope is 1 and the y-intercept is -2.